# Geodesics and Shadows in the Kerr–Bertotti–Robinson Black Hole Spacetime

### v1

@@ 5 School of Physics, Peking University, & Center for High Energy Physics,
$\ast$ Corresponding author: minyongguo@bnu.edu.cn
@@ 1 Introduction
In contrast to Kerr–Melvin black holes, which are of Petrov type I, KBR black holes belong to the Petrov type D class [11]. Notably, they feature bounded ergoregions, and their electromagnetic fields remain asymptotically finite and uniform. In principle, test particles can escape to infinity. These properties make KBR black holes a more physically realistic model of a black hole immersed in an external electromagnetic field. The Schwarzschild–Melvin and Kerr–Melvin solutions have been extensively studied, covering topics such as geodesic structure [12, 13, 14], the dynamics of charged particles [15, 16, 17], black hole shadows [18, 19], photon rings, and image formation from luminous accretion disks [20]. It is therefore natural to explore these phenomena in the KBR spacetime. It is also worth noting that magnetic reconnection in the presence of strong magnetic fields can be significant [21, 22], potentially triggering an astrophysically realistic Penrose process within the ergosphere [23, 24, 25, 26, 27], as recently explored for the KBR black hole in [28].
@@
Moreover, given the growing importance of black hole shadow and image studies [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]—particularly in light of the Event Horizon Telescope’s groundbreaking images of the supermassive black holes at the centers of the M87 galaxy and the Milky Way [59, 60, 61, 62]—the KBR black hole, as a novel class of exact solutions, merits dedicated investigation. In this work, we focus on the analysis of geodesics and the shadow of the KBR black hole, thereby laying the foundation for more comprehensive investigations of its observational features. We show that null geodesics remain analytically tractable due to the separability of the Hamilton–Jacobi equation, whereas timelike geodesics are generally non-separable. We derive approximate analytic expressions for the photon sphere and the innermost stable circular orbit (ISCO), expanded perturbatively in the magnetic field strength. The black hole shadow is then computed using both analytical approximations and numerical ray-tracing techniques. To quantify deviations from the Kerr shadow, we introduce a deformation parameter and analyze its dependence on the magnetic field strength, the observer’s inclination angle, and radial position.
@@
The remaining sections of this paper are structured as follows. In Sec. 2, we review the KBR black hole spacetime and discuss its basic properties. Sec. 3 presents a comprehensive and systematic investigation of the null and timelike geodesics within the KBR black hole spacetime. In Sec. 4, we explore the shadow cast by the KBR black hole and quantitatively examine its deviations from, and distinctions relative to, that of the Kerr black hole. Finally, in Sec. 5, we provide a summary and outlook. In this work, we have set the fundamental constants $c$, $G$ to unity, and we will work in the convention (,+,+,+)(-,+,+,+)( - , + , + , + ).
@@ 2 The KBR Spacetime
Here, $m$ denotes the mass parameter of the central black hole, $a$ is the spin parameter, and $B$ characterizes the strength of the electromagnetic field. The spacetime given by Eq. (2.1) satisfies the Einstein–Maxwell equations, where the 1-form gauge potential is given by:
@@
[…]
@@
From this expression, we observe that when the magnetic field is nonzero, the maximum value of $a$ is less than $m$. This implies that, in the presence of a non-vanishing magnetic field, the spin parameter $a$ of an extremal KBR black hole is strictly less than its mass parameter. In the absence of the Maxwell field, namely when $B=0$, the metric reduces to the Kerr solution expressed in Boyer–Lindquist coordinates. In the limiting case $m\to 0$, the geometry approaches the BR metric, which exhibits the symmetry of a direct product spacetime, $\text{AdS}_{2}\times S^{2}$. When both $B$ and $M$ vanish, the metric degenerates into flat Minkowski spacetime. If only the rotation parameter $a$ is set to zero, the spacetime reduces to the Schwarzschild–BR geometry. The parameters $B$ and $a$ possess dimensions of inverse length and length, respectively:
$[B]=[a^{-1}]=[m^{-1}]=[r^{-1}].$
@@
$B_{\text{Gauss}}=\frac{c^{4}}{G^{3/2}m}B=2.36\times 10^{19}\mathopen{}\mathclose{{\left(\frac{M_{\odot}}{m}}}\right)(Bm),$
@@ 3 Geodesics of the KBR Exterior
where $\mu=1$ for massive particles and $\mu=0$ for massless particles. The four-momentum of the particle is defined by $p_{\mu}=\partial S/\partial x^{\mu}$. In a stationary and axisymmetric spacetime, the existence of two Killing vectors leads to the conservation of energy and angular momentum, expressed as $\mathcal{E}=-p_{t}$, $\mathcal{L}=p_{\phi}$, respectively. The Hamiltonian can be rewritten in the form $2H=\mathcal{C}+g^{rr}p_{r}+g^{\theta\theta}p_{\theta}$, where the function $\mathcal{C}$ is defined as $\mathcal{C}=g^{tt}\mathcal{E}^{2}-2g^{t\phi}\mathcal{E}\mathcal{L}+g^{\phi\phi}\mathcal{L}^{2}$. Substituting the specific form of the KBR metric from Eq. (2.10) into $\mathcal{C}$ , we obtain
$\frac{\mathcal{C}}{\Omega^{2}}$ $=$ $\frac{C^{2}Qa\sin^{2}{\theta}\mathcal{E}^{2}+Q\csc^{2}{\theta}\mathcal{L}^{2}-Pa^{2}\mathcal{L}^{2}-C^{2}P\mathopen{}\mathclose{{\left(r^{2}+a^{2}}}\right)^{2}-2aCQ\mathcal{E}\mathcal{L}+2aCP\mathopen{}\mathclose{{\left(r^{2}+a^{2}}}\right)\mathcal{E}\mathcal{L}}{C^{2}PQ\Sigma}$ (3.3)
@@
Although the terms within the square brackets are separable in $r$ and $\theta$, the prefactor $\Sigma^{-1}\Omega^{2}$, which depends on both variables, generally prevents full separation of variables in the Hamilton–Jacobi equation for timelike particles ($\mu\neq 0$). Consequently, the geodesic equations are, in general, not separable for massive particles. However, for null geodesics ($\mu=0$), the equation becomes separable. In this case, the Hamilton principal function can be written as $S=S_{r}(r)+S_{\theta}(\theta)-\mathcal{E}t+\mathcal{L}\phi$. This separability arises from the fact that the conformal factor does not affect the trajectories of null geodesics. Finally, we note that the parameter $C$ can be absorbed into the angular momentum by the rescaling $\mathcal{L}\rightarrow C\mathcal{L}$ in Eq. (3.4). Without loss of generality, we set $C=1$ in the remainder of this analysis.
@@ 3.1 Null geodesics and photon region
In this subsection, we focus on null geodesics. In this case, the radial and angular components of the four-momentum are given by $p_{r}=S_{r}(r)$ and $p_{\theta}=S_{\theta}(\theta)$, respectively. Setting $\mu=0$ in Eq. (3.4), the equations of motion in the $r$ and $\theta$ directions become
@@
where $\kappa$ is a separation constant arising from the separability of the Hamilton–Jacobi equation. This constant is independent of the particle’s mass, energy, or angular momentum. It is convenient to define a shifted conserved quantity,
@@
$\frac{\Sigma}{\Omega^{2}}\frac{\mathrm{d}r}{\mathrm{d}\lambda}$ =±$=\pm_{r}\sqrt{R(r)},$
$\frac{\Sigma}{\Omega^{2}}\frac{\mathrm{d}\theta}{\mathrm{d}\lambda}$ =±$=\pm_{\theta}\sqrt{\Theta(\theta)},$
@@
To compute null geodesics, we first solve the lensing equation, which is obtained from the ratio of the radial and angular components in Eq. (3.9), yielding a relation between the variations in $r$ and $\theta$. Once this relation is established, the equations for $t,\phi$ can be integrated to determine the corresponding changes in coordinate time and azimuthal angle.
@@
$A_{1}$ $=\mathopen{}\mathclose{{\left[a^{2}-\lambda^{2}-\eta-B^{2}\mathopen{}\mathclose{{\left(a^{2}-\frac{I_{2}}{I_{1}^{2}}}}\right)\mathopen{}\mathclose{{\left[(a-\lambda)^{2}+\eta}}\right]}}\right]\,,$ (3.12)
$A_{2}$ $=\mathopen{}\mathclose{{\left[a^{2}-B^{2}\mathopen{}\mathclose{{\left(a^{2}-\frac{I_{2}}{I_{1}^{2}}}}\right)\mathopen{}\mathclose{{\left[(a-\lambda)^{2}+\eta}}\right]}}\right]\,,\quad u_{\pm}=\frac{A_{1}\pm\sqrt{A_{1}^{2}+4\eta}}{2A_{2}}\,.$
For small values of $B$, and assuming $a\neq 0$, we have $A_{2}>0$, implying that the qualitative behavior of motion in the $\theta$ direction remains similar to that in the Kerr family spacetimes [66, 67]. Since $\Theta|_{u\to 1}\to-\lambda\mathopen{}\mathclose{{\left(1-u}}\right)^{-1}$, only photons with $\lambda=0$ can reach the polar axis. When $\eta\geq 0$, one finds $u_{-}<0$, meaning that $u\in[0,u_{+}]$, and photons can cross the equatorial plane ($u=0$), oscillating within the range $\theta\in[\theta_{+},\pi-\theta_{+}]$. Conversely, for $\eta<0$, we have $u_{-}>0$, and photons are confined to the regions $\theta\in[\theta_{-},\theta_{+}]$ or $[\pi-\theta_{+},\pi-\theta_{-}]$, where $)\theta_{\pm}=\arccos{\sqrt{u_{\pm}}}$.
@@ 3.2 Circular timelike geodesics and the ISCO
$\Omega_{(B)}=\frac{-4a^{5}+2a^{3}\mathopen{}\mathclose{{\left(r^{3}+2}}\right)-4ar^{3}\mathopen{}\mathclose{{\left(1+2r^{2}+r^{3}}}\right)\pm 3a^{4}r^{3/2}\pm 4r^{9/2}\mathopen{}\mathclose{{\left(r^{2}+1}}\right)\pm a^{2}r^{3/2}\mathopen{}\mathclose{{\left(3r^{3}+4r^{2}-4}}\right)}{4\mathopen{}\mathclose{{\left(r^{3}-a^{2}}}\right)^{2}}\,.$ (3.20)
@@ 4.1 Set up of the observer
[…]
@@ 4.2 The critical curve
In this subsection, we present the critical curve structures in the KBR spacetime. Figure 2 illustrates how the critical curves vary with different black hole parameters and observer positions. The top-left panel shows the case of a relatively high spin parameter, $a=0.94$, with the observer located at a small inclination angle $\theta_{0}=17^{\circ}$. In this case, the critical curve is nearly circular. Notably, the size of the shadow is sensitive to the magnetic field strength $B$: inceasing $B$ enlarges the critical curve, while its circularity remains largely unaffected. A similar trend is observed in the top-right panel, where the observer is placed at the equatorial plane ($\theta_{0}=90^{\circ}$). However, due to the black hole’s rotation, the critical curve exhibits a characteristic “D”-shaped distortion.
@@
Applying this analysis to the case shown in Figure 2, the characteristic radius takes the values of 100100100 for $B=0.01$ and 200200200 for $B=0.005$. Accordingly, an observer located at $r_{o}=300$ effectively resides in the far zone, where the asymptotic deviations from the Kerr geometry become significant. This observer is thus capable of detecting the modifications to the black hole shadow induced by the magnetic field. In contrast, an observer situated at $r_{o}=50$ effectively lies well within the near zone, where the spacetime remains effectively Kerr-like and the asymptotically AdS-like structure is not observable.
@@ Acknowledgments
We are grateful to Jinchi Liu for his valuable support. The work is partly supported by NSFC Grant No. 12205013 and 12275004.

### v2

@@ 5 School of Physics, Peking University, & Center for High Energy Physics,
$\ast$ Corresponding authors: minyongguo@bnu.edu.cn, chenbin1@nbu.edu.cn
@@ 1 Introduction
In contrast to Kerr–Melvin black holes, which are of Petrov type I [8], KBR black holes belong to the Petrov type D class [62]. Notably, they feature bounded ergoregions, and their electromagnetic fields remain asymptotically finite and uniform. In principle, test particles can escape to infinity. These properties make KBR black holes a more physically realistic model of a black hole immersed in an external electromagnetic field. The Schwarzschild–Melvin and Kerr–Melvin solutions have been extensively studied, covering topics such as geodesic structure [66, 21, 25], the dynamics of charged particles [19, 51, 64], black hole shadows [45, 69], photon rings, and image formation from luminous accretion disks [39]. It is therefore natural to explore these phenomena in the KBR spacetime. It is also worth noting that magnetic reconnection in the presence of strong magnetic fields can be significant [26, 27], potentially triggering an astrophysically realistic Penrose process within the ergosphere [48, 17, 14, 78, 65], as recently explored for the KBR black hole in [75].
@@
Moreover, given the growing importance of black hole shadow and image studies [29, 46, 13, 34, 35, 82, 37, 57, 56, 77, 76, 49, 31, 81, 40, 74, 36, 58, 20, 32, 53, 54, 59, 42, 52, 55, 70, 72, 73, 33, 15, 43, 71, 50, 67]—particularly in light of the Event Horizon Telescope’s groundbreaking images of the supermassive black holes at the centers of the M87 galaxy and the Milky Way [2, 5, 3, 4]—the KBR black hole, as a novel class of exact solutions, merits dedicated investigation. In this work, we focus on the analysis of geodesics and the shadow of the KBR black hole, thereby laying the foundation for more comprehensive investigations of its observational features. We show that null geodesics remain analytically tractable due to the separability of the Hamilton–Jacobi equation, whereas timelike geodesics are generally non-separable. We derive approximate analytic expressions for the photon sphere and the innermost stable circular orbit (ISCO), expanded perturbatively in the magnetic field strength. The black hole shadow is then computed using both analytical approximations and numerical ray-tracing techniques. To quantify deviations from the Kerr shadow, we introduce a deformation parameter and analyze its dependence on the magnetic field strength, the observer’s inclination angle, and radial position.
@@
The remaining sections of this paper are structured as follows. In Sec. 2, we review the KBR black hole spacetime and discuss its basic properties. Sec. 3 presents a comprehensive and systematic investigation of the null and timelike geodesics within the KBR black hole spacetime. In Sec. 4, we explore the shadow cast by the KBR black hole and quantitatively examine its deviations from, and distinctions relative to, that of the Kerr black hole. Finally, in Sec. 5, we provide a summary and outlook. In this work, we have set the fundamental constants $c$, $G$ to unity, and we will work in the convention (,+,+,+)(-,+,+,+).
@@ 2 The KBR Spacetime
Here, $m$ denotes the mass parameter of the central black hole, $a$ is the spin parameter, and $B$ characterizes the strength of the electromagnetic field. The spacetime given by Eq. (2.1) satisfies the Einstein–Maxwell equations, where the 1-form gauge potential is given by
@@
$2\pi$ $=\lim_{\theta\to 0}\frac{\text{circumference}}{\text{radius}}=\int_{0}^{\varphi_{\text{max}}}\sqrt{g_{\varphi\varphi}}\,d\varphi=\varphi_{\text{max}}P(0)\,$,
$2\pi$ $=\lim_{\theta\to\pi}\frac{\text{circumference}}{\text{radius}}=\int_{0}^{\varphi_{\text{max}}}\sqrt{g_{\varphi\varphi}}\,d\varphi=\varphi_{\text{max}}P(\pi)\,.$ (2.9)
@@
From this expression, we observe that when the magnetic field is nonzero, the maximum value of $a$ is less than $m$. This implies that, in the presence of a non-vanishing magnetic field, the spin parameter $a$ of an extremal KBR black hole is strictly less than its mass parameter. In the absence of the Maxwell field, namely when $B=0$, the metric reduces to the Kerr solution expressed in Boyer–Lindquist coordinates. In the limiting case $m\to 0$, the geometry approaches the BR metric [6], which exhibits the symmetry of a direct product spacetime, $\text{AdS}_{2}\times S^{2}$. When both $B$ and $m$ vanish, the metric degenerates into flat Minkowski spacetime. If only the rotation parameter $a$ is set to zero, the spacetime reduces to the Schwarzschild–BR geometry. The parameters $B$ and $a$ possess dimensions of inverse length and length, respectively:
$B]=[a^{-1}]=[m^{-1}]=[r^{-1}].$
@@
$B_{\text{Gauss}}=\frac{c^{4}}{G^{3/2}m}(Bm)=2.36\times 10^{19}\mathopen{}\mathclose{{\left(\frac{M_{\odot}}{m}}}\right)(Bm),$
@@ 3 Geodesics of the KBR Exterior
where $\mu=1$ for massive particles and $\mu=0$ for massless particles. The four-momentum of the particle is defined by $p_{\mu}=\partial S/\partial x^{\mu}$. In a stationary and axisymmetric spacetime, the existence of two Killing vectors leads to the conservation of energy and angular momentum, expressed as $\mathcal{E}=-p_{t}$, $\mathcal{L}=p_{\phi}$, respectively. The Hamiltonian can be rewritten in the form $2H=\mathcal{C}+g^{rr}p_{r}^{2}+g^{\theta\theta}p_{\theta}^{2}$, where the function $\mathcal{C}$ is defined as $\mathcal{C}=g^{tt}\mathcal{E}^{2}-2g^{t\phi}\mathcal{E}\mathcal{L}+g^{\phi\phi}\mathcal{L}^{2}$. Substituting the specific form of the KBR metric from Eq. (2.10) into $\mathcal{C}$, we obtain
$\frac{\mathcal{C}}{\Omega^{2}}$ $=$ $\frac{C^{2}Qa\sin^{2}{\theta}\mathcal{E}^{2}+Q\csc^{2}{\theta}\mathcal{L}^{2}-Pa^{2}\mathcal{L}^{2}-C^{2}P\mathopen{}\mathclose{{\left(r^{2}+a^{2}}}\right)^{2}\mathcal{E}^{2}-2aCQ\mathcal{E}\mathcal{L}+2aCP\mathopen{}\mathclose{{\left(r^{2}+a^{2}}}\right)\mathcal{E}\mathcal{L}}{C^{2}PQ\Sigma}$ (3.3)
@@
Although the terms within the square brackets are separable in $r$ and $\theta$, the prefactor $\Sigma^{-1}\Omega^{2}$, which depends on both variables, generally prevents full separation of variables in the Hamilton–Jacobi equation for timelike particles ($\mu\neq 0$). Consequently, the geodesic equations are, in general, not separable for massive particles. However, for null geodesics ($\mu=0$), the equation becomes separable. In this case, the Hamilton principal function can be written as $S=S_{r}(r)+S_{\theta}(\theta)-\mathcal{E}t+\mathcal{L}\phi$. This separability arises from the fact that the conformal factor does not affect the trajectories of null geodesics. Finally, we note that the parameter $C$ can be absorbed into the angular momentum by the rescaling $\mathcal{L}\rightarrow C\mathcal{L}$ in Eq. (3.4). Without loss of generality, we set $C=1$ for the remainder of this analysis.
@@ 3.1 Null geodesics and photon region
In this subsection, we focus on null geodesics. In this case, the radial and angular components of the four-momentum are given by $p_{r}=\partial_{r}S_{r}(r)$ and $p_{\theta}=\partial_{\theta}S_{\theta}(\theta)$, respectively. Setting $\mu=0$ in Eq. (3.4), the equations of motion in the $r$ and $\theta$ directions become
@@
where $\kappa$ is a separation constant arising from the separability of the Hamilton–Jacobi equation. This constant is independent of the particle’s mass, energy, and angular momentum. It is convenient to define a shifted conserved quantity,
@@
$\frac{\Sigma}{\Omega^{2}}\frac{\mathrm{d}r}{\mathrm{d}\lambda}$ $=\pm_{r}\sqrt{R(r)}\,$,
$\frac{\Sigma}{\Omega^{2}}\frac{\mathrm{d}\theta}{\mathrm{d}\lambda}$ $=\pm_{\theta}\sqrt{\Theta(\theta)}\,$,
@@
To compute null geodesics, we first solve the lensing equation, which is obtained from the ratio of the radial and angular components in Eq. (3.9), yielding a relation between the variations in $r$ and $\theta$. Once this relation is established, the equations for $t$ and $\phi$ can be integrated to determine the corresponding changes in coordinate time and azimuthal angle.
@@
$A_{1}$ $=\mathopen{}\mathclose{{\left[a^{2}-\xi^{2}-\eta-B^{2}\mathopen{}\mathclose{{\left(a^{2}-\frac{I_{2}}{I_{1}^{2}}}}\right)\mathopen{}\mathclose{{\left[(a-\xi)^{2}+\eta}}\right]}}\right]\,$, (3.12)
$A_{2}$ $=\mathopen{}\mathclose{{\left[a^{2}-B^{2}\mathopen{}\mathclose{{\left(a^{2}-\frac{I_{2}}{I_{1}^{2}}}}\right)\mathopen{}\mathclose{{\left[(a-\xi)^{2}+\eta}}\right]}}\right]\,,\quad u_{\pm}=\frac{A_{1}\pm\sqrt{A_{1}^{2}+4\eta A_{2}}}{2A_{2}}\,.$
[…]
@@ 3.2 Circular timelike geodesics and the ISCO
$\Omega_{(B)}=\frac{r^{3/2}\mathopen{}\mathclose{{\left(a^{2}\pm 4ar^{3/2}-4r^{2}}}\right)}{4\mathopen{}\mathclose{{\left(a\pm r^{3/2}}}\right)^{2}}\,.$ (3.20)
@@ 4.1 Set up of the observer
It is worth noting that, under the definition, the range of $\Psi$ spans from 0 to $\pi$. The corresponding Cartesian coordinates $(x,y)$ on the image plane are then obtained via the stereographic projection formulas:
@@ 4.2 The critical curve
In this subsection, we present the critical curve structures in the KBR spacetime. Figure 2 illustrates how the critical curves vary with different black hole parameters and observer positions. The top-left panel shows the case of a relatively high spin parameter, $a=0.94$, with the observer located at a small inclination angle $\theta_{o}=17^{\circ}$. In this case, the critical curve is nearly circular. Notably, the size of the shadow is sensitive to the magnetic field strength $B$: increasing $B$ enlarges the critical curve, while its circularity remains largely unaffected. A similar trend is observed in the top-right panel, where the observer is placed at the equatorial plane ($\theta_{o}=90^{\circ}$). However, due to the black hole’s rotation, the critical curve exhibits a characteristic “D”-shaped distortion.
@@
Applying this analysis to the case shown in Figure 2, the characteristic radius takes the values of $100$ for $B=0.01$ and $200$ for $B=0.005$. Accordingly, an observer located at $r_{o}=300$ effectively resides in the far zone, where the asymptotic deviations from the Kerr geometry become significant. This observer is thus capable of detecting the modifications to the black hole shadow induced by the magnetic field. In contrast, an observer situated at $r_{o}=50$ effectively lies well within the near zone, where the spacetime remains effectively Kerr-like and the asymptotically AdS-like structure is not observable.
@@ Acknowledgments
We are grateful to Jinchi Liu for his valuable support. The work is partly supported by NSFC Grant No. 12205013, 12275004 and 12588101. Y. H. acknowledges support from the NSFC Grant No. 12547123. M. G. is also supported by Open Fund of Key Laboratory of Multiscale Spin Physics (Ministry of Education), Beijing Normal University and he is also supported by the BNU Tang Scholar.